# Probing the nature of gravity in the low-acceleration limit: wide binaries of extreme separations with perspective effects

### v1

@@ Abstract
Recent statistical analyses of wide binaries have revealed a boost in gravitational acceleration with respect to the prediction by Newtonian gravity at low internal accelerations $\lesssim 10^{-9}$ m s-2. This phenomenon is important because it does not permit the dark matter interpretation, unlike galaxy rotation curves. We extend previous analyses by increasing the maximum sky-projected separation from 30 to 50 kilo astronomical units (kau). We show that the so-called “perspective effects” are not negligible at this extended separation and, thus, incorporate it in our analysis. With wide binaries selected with very stringent criteria, we find that the gravitational acceleration boost factor, $\gamma_{g}\equiv g_{\rm obs}/g_{\mathrm{N}}$, is $1.61^{+0.37}_{-0.29}$ (from $\delta_{\rm obs-newt}\equiv(\log_{10}\gamma_{g})/\sqrt{2}=0.147\pm 0.062$) at Newtonian accelerations $g_{\mathrm{N}}=10^{-11.0}$ m s-2, corresponding to separations of tens of kau for solar-mass binaries. At Newtonian accelerations $g_{\mathrm{N}}=10^{-10.3}$ m s-2, we find $\gamma_{g}=1.26^{+0.12}_{-0.10}$ ($\delta_{\rm obs-newt}=0.072\pm 0.027$). For all binaries with $g_{\rm N}\lesssim 10^{-10}$ m $s^{-2}$ from our sample, we find $\gamma_{g}=1.32^{+0.12}_{-0.11}$ ($\delta_{\rm obs-newt}=0.085\pm 0.027$). These results are consistent with the generic prediction of MOND-type modified gravity, although the current data are not sufficient to pin down the low-acceleration limiting behavior. Finally, we emphasize that the observed deviation from Newtonian gravity cannot be explained by the perspective effects or any separation-dependent eccentricity variation which we take into account.
@@ 1 Introduction
In 1983, Milgrom proposed Modified Newtonian Dynamics (MOND, M. Milgrom, 1983) as an alternative path to the dark matter problem. For MOND reviews, the reader is referred to R. H. Sanders & S. S. McGaugh (2002), B. Famaey & S. S. McGaugh (2012), and I. Banik & H. Zhao (2022). MOND posits an acceleration scale $a_{0}$ ($\approx 1.2\times 10^{-10}$ m s-2) as a new physical constant that marks the scale where standard gravity breaks down. This basic tenet of MOND can be most directly tested by wide binaries because any hypothetical dark matter cannot play any role in their internal dynamics (X. Hernandez et al., 2012). When the two stars of a binary are separated by more than several kau, their mutual internal acceleration gets $\lesssim a_{0}$ and, according to MOND, the norm of the relative velocity between the pair must be greater than the Newtonian expectation. On the other hand, dark matter can hardly change the relative velocity, given that its hypothetical total mass within a wide binary orbit is smaller than the wide binary mass by more than four orders of magnitude. Thus, if wide binary orbit anomalies are experimentally proven, it means a falsification of the assumption of standard gravity behind dark matter.
@@
By analyzing wide binaries from the Gaia data release (DR) 3 database (Gaia Collaboration et al., 2023), Chae (K.-H. Chae, 2023, 2024a, 2024b, 2025) and Hernandez et al. (X. Hernandez, 2023; X. Hernandez et al., 2024; X. Hernandez & P. Kroupa, 2025) independently concluded that gravitational acceleration is boosted by a factor of $\gamma_{g}\approx 1.4-1.5$ at Newton-predicted gravitational acceleration $g_{\rm N}\lesssim 10^{-9.5}$ m s-2 in agreement with the generic predictions of Lagrangian theories of nonrelativistic MOND gravity such as AQUAL (J. Bekenstein & M. Milgrom, 1984) and QUMOND (M. Milgrom, 2010) and relativistic MOND gravity (see e.g., T. Mistele, 2024). These theories break the strong equivalence principle (while keeping Galilei’s universality of free fall), necessitating the external field effect (EFE). According to such theories, wide binaries in the solar neighborhood are subject to an EFE from the rather strong external field $\approx 1.8a_{0}$ from the Milky Way and are consequently expected to obey a rather mild gravity boost of $\approx 1.4$.
@@
However, based on the currently available data, there are two major observational limitations in using wide binaries to check whether they deviate from Newtonian predictions. First, Newtonian gravity predicts an elliptical orbit for a bound binary system, and a significant segment of the orbit is ideally needed to determine the orbital shape, but the DR3 time span is 34 months, while the orbital periods of typical wide binaries are on the order of hundreds of thousands of years. Second, even from the “snapshot” observations of wide binaries, not all of the six components of the relative displacement and relative velocity between the pair are accurately measured. In general, the line-of-sight separation is essentially unknown or ill-determined (as the distance errors are usually much larger than the relatively small orbital sizes although they are “wide” binaries), and radial velocities of the stars, in general, have large uncertainties (or are not available at all) from the Gaia DR3. Therefore, most studies used statistical methods based only on the 2D relative position and the 2D relative velocity projected on the sky plane, which are in general accurately measured from Gaia observations (see Table 3 of K.-H. Chae (2025) for a summary of statistical methods).
@@
Recently, new approaches beyond statistical methods relying only on 2D quantities have been proposed or considered. K.-H. Chae (2025) considered a sample of 312 wide binaries selected from Gaia DR3 that have relatively accurate radial velocities, while R. Saglia et al. (2025) considered a smaller sample of 32 binaries (including only several with internal acceleration $\lesssim a_{0}$) that have the HARPS radial velocities whose precision is typically tens of m s-1 (an order of magnitude more precise than the currently available Gaia radial velocities). K.-H. Chae (2025) also employed a novel Bayesian method to derive a probability distribution of the gravity boost factor. As regards these new developments, R. Saglia et al. (2025) is limited by the small sample size while K.-H. Chae (2025) is limited by the relatively larger uncertainties of radial velocities, and both are limited by insufficiently precise distances. However, these approaches will become more and more fruitful as more and more sufficiently precise radial velocities become available in the future.
@@
Regardless of methods to probe gravity, wide binaries with extremely large separations have not yet been analyzed properly, although they are potentially of great importance because they can be used to probe the low-acceleration limiting behavior of gravity. Use of extremely wide binaries involves additional difficulties. First, it is more difficult to find truly isolated and gravitationally-bound wide binaries as the separation gets larger. Thus, extreme care is needed to select those binaries. Second, the relative velocity between the pair gets smaller at larger separation, and thus greater measurement precision is required. Finally, the so-called “perspective effects” (E. Shaya & R. Olling, 2011) are important and have to be taken into account. When measuring the relative velocity, measured velocities can simply be subtracted from each other for binaries with small separations, as their relative positions can be regarded as being placed on a flat surface due to the negligible curvature of the celestial sphere. However, this is not the case for those with large separations, especially if the barycenter of the binary system has a large velocity relative to the Sun.
@@
The organization of this paper is as follows. In Section 2, we describe our selection criteria for binaries. In Section 3, we describe how we calculate the perspective effects. In Section 4, we describe some scaling relations in our sample. In Section 5, we present our Monte Carlo (MC) simulation results. Discussions and conclusions are given in Section 6. The newly defined wide sample will be made public at Zenodo. The codes used for statistical analyses can be found at K.-H. Chae (2024c).
@@ 2 Wide Binaries Sample
Our analysis is based on the comprehensive catalog of wide binaries constructed by K. El-Badry et al. (2021) from the Gaia DR3 database. This catalog is based on astrometric and photometric quality cuts, and includes pairs with projected separations from a few au up to 1 pc ($2\times 10^{5}$ au) and distances within 1 kpc from the Sun. Importantly, K. El-Badry et al. (2021) empirically estimated the probability of chance alignment $R$ for each pair, providing a reasonable means to isolate gravitationally bound binaries. Their catalog contains approximately $1.3$ million binaries with $>90\%$ probability of being gravitationally bound, and over $1.1$ million with a $>99\%$ probability, making it the largest and most reliable sample of wide binaries currently available. Bound binaries in this context are defined as pairs whose kinematics and parallaxes are consistent, within uncertainties, with being gravitationally bound.
@@
Figure 1: Color–magnitude diagrams for the primary (component A, left panel) and secondary (component B, right panel) stars in our wide binary sample. Each point represents a star, plotted in absolute $G$-band magnitude ($M_{G}$) versus Gaia color ($BP-RP$). The shaded blue region denotes the selection window applied to ensure that both components are main-sequence stars and to exclude evolved objects and outliers. The dashed red lines indicate the magnitude and color boundaries: $13.43.8<M_{G}<$ 13.4 and $M_{G}-3.2(BP-RP)<3.8$. Only binaries in which both components fall within the shaded region are retained for subsequent analysis.
@@
  • As in K.-H. Chae (2023), we account for dust extinction using a 3D dust map available only for declinations above $-28^{\circ}$; therefore, binaries located below this declination are excluded, removing about 30% of the data.
@@ 3 Perspective effects
We note that the analytically calculated perspective effects given by Equations (10), (11), and (12) agree well with direct numerical calculations with the algorithm described in Appendix B of K.-H. Chae (2025).
@@
is about two orders of magnitude greater than the true value of $\Delta d$, which should be in the order of $s$, the sky-projected separation. See the top-left panel of Figure 2 for a histogram of $\sigma_{\Delta d}/s$. Thus, considering the first term would induce errors that would be much larger than those without it. Lacking accurate values of $\Delta d$, we could estimate the effect of the term proportional to $\Delta d/d$ using realistic MC generated values of $\Delta d$, but it is not warranted in a statistical sense because $\Delta d/d$ can be positive or negative with equal probabilities, see Appendix B of K.-H. Chae (2025).
@@ 4 Scaling relations
Then $s/r_{M}$ (where $r_{M}$ is given by Equation (18)) and $\tilde{v}$ are dimensionless. According to Newtonian gravity, $\tilde{v}$ should be roughly constant, independent of $s/r_{M}$ (as long as eccentricity is nearly independent of $s$). However, we clearly see that $\tilde{v}$ deviates from a flat line beyond about $s/r_{M}=0.3$.
@@ 5 Monte Carlo simulations
For the present sample, the two highest acceleration bins with $x_{0}>-8$ satisfy $\delta_{\rm obs-newt}\approx 0$ with $f_{\rm multi}=0$. This means that in an average or median sense, our sample is pure, free of kinematic contaminants. This can be understood from the fact that every binary in our sample satisfies the highest data qualities (e.g., $\sigma_{v_{p}}<50$ m s-1) and stringent selection criteria such as ruwe $<1.2$. This will not mean that there are no individual exceptions that evaded our selection criteria, but relatively few exceptions would not affect the median trend. We also emphasize that every star from all our binaries satisfies the same criteria with an equal probability, regardless of their membership, meaning that the exceptional cases would occur regardless of the binary separation and thus cannot bias the median trend we see.
The purple line in Figure 6 is the AQUAL prediction, as presented in Figure 1 of K.-H. Chae (2024a). We find that the gravitational acceleration boost factor $\gamma_{g}$ (Equation (24)) is $1.61^{+0.37}_{-0.29}$ in the last bin, i.e., at $g_{\rm N}\approx 10^{-11.0}\text{m s}^{-2}$ and $1.26^{+0.12}_{-0.10}$ in the second to last bin, i.e., at $g_{\rm N}\approx 10^{-10.3}\text{m s}^{-2}$. The last bin represents the lowest-acceleration bin probed to date, thanks to our inclusion of wide binaries with $s>30$ kau. Unfortunately, it returns a result with a relatively large uncertainty. Nevertheless, it is intriguing to note that the boost factor $1.61^{+0.37}_{-0.29}$ is consistent with the AQUAL/QUMOND prediction $\approx 1.4-1.5$. Although not shown in the figure, we find $1.32^{+0.12}_{-0.11}$ for $g_{\rm N}\lesssim 10^{-10}\text{m s}^{-2}$ that encompasses the two bins. These boost factors are consistent with previous estimates based on differently defined samples (e.g., K.-H. Chae 2023, 2024a, 2024b, 2025; X. Hernandez et al. 2024; X. Hernandez & P. Kroupa 2025).
@@
The above results are based on the nominal sample and the 2D velocities corrected for the perspective effects. In Appendix A, we perform the acceleration-plane analysis of our nominal sample without considering the perspective effects, so that the impact of the perspective effects on gravity inference can be seen directly. We see that the perspective effects have a rather minor, though non-negligible, impact on gravity inference. In Appendix B, we consider alternative samples with varied ruwe limits of $<1.1$ and $<1.3$ instead of the nominal limit $<1.2$. We see that there are no significant changes in the gravity inference.

### v2

@@ Abstract
Recent statistical analyses of wide binaries have revealed a boost in gravitational acceleration with respect to the prediction by Newtonian gravity at low internal accelerations $\lesssim 10^{-9}$ m s-2. This phenomenon is important because it does not permit the dark matter interpretation, unlike galaxy rotation curves. We extend previous analyses by increasing the maximum sky-projected separation from 30 to 50 kilo astronomical units (kau). We show that the so-called “perspective effects” are not negligible at this extended separation and, thus, incorporate it in our analysis. With wide binaries selected with very stringent criteria, we find that the gravitational acceleration boost factor, $\gamma_{g}\equiv g_{\rm obs}/g_{\mathrm{N}}$, is $1.61^{+0.37}_{-0.29}$ (from $\delta_{\rm obs-newt}\equiv(\log_{10}\gamma_{g})/\sqrt{2}=0.147\pm 0.062$) at Newtonian accelerations $g_{\mathrm{N}}=10^{-11.0}$ m s-2, corresponding to separations of tens of kau for solar-mass binaries. At Newtonian accelerations $g_{\mathrm{N}}=10^{-10.3}$ m s-2, we find $\gamma_{g}=1.26^{+0.12}_{-0.10}$ ($\delta_{\rm obs-newt}=0.072\pm 0.027$). For all binaries with $g_{\rm N}\lesssim 10^{-10}$ m $s^{-2}$ from our sample, we find $\gamma_{g}=1.32^{+0.12}_{-0.11}$ ($\delta_{\rm obs-newt}=0.085\pm 0.027$). These results are consistent with the generic prediction of MOND-type modified gravity, although the current data are not sufficient to pin down the low-acceleration limiting behavior. Finally, we emphasize that the observed deviation from Newtonian gravity cannot be explained by the perspective effects or any separation-dependent eccentricity variation that we have taken into account.
@@ 1 Introduction
In 1983, Milgrom proposed Modified Newtonian Dynamics (MOND, M. Milgrom, 1983) as an alternative path to the dark matter problem. For MOND reviews, the reader is referred to R. H. Sanders & S. S. McGaugh (2002), B. Famaey & S. S. McGaugh (2012), and I. Banik & H. Zhao (2022). MOND posits an acceleration scale $a_{0}$ ($\approx 1.2\times 10^{-10}$ m s-2) as a new physical constant that marks the scale where standard gravity breaks down. This basic tenet of MOND can be most directly tested by wide binaries because any hypothetical dark matter cannot play any role in their internal dynamics (X. Hernandez et al., 2012). When the two stars of a binary are separated by more than several kau, their mutual internal acceleration gets $\lesssim a_{0}$ and, according to MOND, the norm of the relative velocity between the pair must be greater than the Newtonian expectation. On the other hand, dark matter can hardly change the relative velocity, given that its hypothetical total mass within a wide binary orbit is smaller than the wide binary mass by more than four orders of magnitude. The dark matter density near the Sun is estimated to be only $10^{-9}M_{\odot}$/kau3 (J. I. Read, 2014). Therefore, the mass of dark matter inside a few kau sphere is less than $10^{-4}M_{\odot}$. Thus, if wide binary orbit anomalies are experimentally proven, it means a falsification of the assumption of standard gravity behind dark matter.
@@
By analyzing wide binaries from the Gaia data release (DR) 3 database (Gaia Collaboration et al., 2023), Chae (K.-H. Chae, 2023, 2024a, 2024b, 2025a) and Hernandez et al. (X. Hernandez, 2023; X. Hernandez et al., 2024; X. Hernandez & P. Kroupa, 2025) independently concluded that gravitational acceleration is boosted by a factor of $\gamma_{g}\approx 1.4-1.5$ at Newton-predicted gravitational acceleration $g_{\rm N}\lesssim 10^{-9.5}$ m s-2 in agreement with the generic predictions of Lagrangian theories of nonrelativistic MOND gravity such as AQUAL (J. Bekenstein & M. Milgrom, 1984) and QUMOND (M. Milgrom, 2010) and relativistic MOND gravity (see e.g., T. Mistele, 2024). These theories break the strong equivalence principle (while keeping Galilei’s universality of free fall), necessitating the external field effect (EFE). According to such theories, wide binaries in the solar neighborhood are subject to an EFE from the rather strong external field $\approx 1.8a_{0}$ from the Milky Way and are consequently expected to obey a rather mild gravity boost of $\approx 1.4$.
@@
However, based on the currently available data, there are two major observational limitations in using wide binaries to check whether they deviate from Newtonian predictions. First, Newtonian gravity predicts an elliptical orbit for a bound binary system, and a significant segment of the orbit is ideally needed to determine the orbital shape, but the DR3 time span is 34 months, while the orbital periods of typical wide binaries are on the order of hundreds of thousands of years. Second, even from the “snapshot” observations of wide binaries, not all of the six components of the relative displacement and relative velocity between the pair are accurately measured. In general, the line-of-sight separation is essentially unknown or ill-determined (as the distance errors are usually much larger than the relatively small orbital sizes although they are “wide” binaries), and radial velocities of the stars, in general, have large uncertainties (or are not available at all) from the Gaia DR3. Therefore, most studies used statistical methods based only on the 2D relative position and the 2D relative velocity projected on the sky plane, which are in general accurately measured from Gaia observations (see Table 3 of K.-H. Chae (2025a) for a summary of statistical methods).
@@
Recently, new approaches beyond statistical methods relying only on 2D quantities have been proposed or considered. K.-H. Chae (2025a) considered a sample of 312 wide binaries selected from Gaia DR3 that have relatively accurate radial velocities, while R. Saglia et al. (2025) considered a smaller sample of 32 binaries (including only several with internal acceleration $\lesssim a_{0}$) that have the HARPS radial velocities whose precision is typically tens of m s-1 (an order of magnitude more precise than the currently available Gaia radial velocities). K.-H. Chae (2025a) also employed a novel Bayesian method to derive a probability distribution of the gravity boost factor. As regards these new developments, R. Saglia et al. (2025) is limited by the small sample size while K.-H. Chae (2025a) is limited by the relatively larger uncertainties of radial velocities, and both are limited by insufficiently precise distances. However, these approaches will become more and more fruitful as more and more sufficiently precise radial velocities become available in the future.
@@
Regardless of methods to probe gravity, wide binaries with extremely large separations have not yet been analyzed properly, although they are potentially of great importance because they can be used to probe the low-acceleration limiting behavior of gravity. Use of extremely wide binaries involves additional difficulties. First, it is more difficult to find truly isolated and gravitationally-bound wide binaries as the separation gets larger. Thus, extreme care is needed to select those binaries. Second, the relative velocity between the pair gets smaller at larger separation, and thus greater measurement precision is required. Finally, the so-called “perspective effects” (E. Shaya & R. Olling, 2011; K. El-Badry, 2019; X. Hernandez, 2023; X. Hernandez et al., 2024) are important and have to be taken into account. When measuring the relative velocity, measured velocities can simply be subtracted from each other for binaries with small separations, as their relative positions can be regarded as being placed on a flat surface due to the negligible curvature of the celestial sphere. However, this is not the case for those with large separations, especially if the barycenter of the binary system has a large velocity relative to the Sun.
@@
The organization of this paper is as follows. In Section 2, we describe our selection criteria for binaries. In Section 3, we describe how we calculate the perspective effects. In Section 4, we describe some scaling relations in our sample. In Section 5, we present our Monte Carlo (MC) simulation results. Discussions and conclusions are given in Section 6. The newly defined wide binary samples are available on Zenodo under an open-source Creative Commons Attribution license: https://doi.org/10.5281/zenodo.16934778 (catalog 10.5281/zenodo.16934778). The codes used for statistical analyses can be found at K.-H. Chae (2024c).
@@ 2 Wide Binaries Sample
Our analysis is based on the comprehensive catalog of wide binaries constructed by K. El-Badry et al. (2021) from the Gaia DR3 database. This catalog is based on astrometric and photometric quality cuts, and includes pairs with projected separations from a few au up to 1 pc ($2\times 10^{5}$ au) and distances within 1 kpc from the Sun. Importantly, K. El-Badry et al. (2021) empirically estimated the probability of chance alignment $R$ for each pair, providing a reasonable means to isolate gravitationally bound binaries. Their catalog contains approximately $1.3$ million binaries with $90\%$ probability of being gravitationally bound, and over $1.1$ million with a $99\%$ probability, making it the largest and most reliable sample of wide binaries currently available. Bound binaries in this context are defined as pairs whose kinematics and parallaxes are consistent, within uncertainties, with being gravitationally bound.
@@
Figure 1: Color–magnitude diagrams for the primary (component A, left panel) and secondary (component B, right panel) stars in our wide binary sample. Each point represents a star, plotted in absolute $G$-band magnitude ($M_{G}$) versus Gaia color (BPRPBP-RP). The shaded blue region denotes the selection window applied to ensure that both components are main-sequence stars and to exclude evolved objects and outliers. The dashed red lines indicate the magnitude and color boundaries: $3.8<M_{G}<13.4$ and $M_{G}-3.2(BP-RP)<3.8$. Only binaries in which both components fall within the shaded region are retained for subsequent analysis.
@@
  • As in K.-H. Chae (2023), we account for dust extinction using a 3D dust map available only for declinations above $28^{\circ}$; therefore, binaries located below this declination are excluded, removing about 30% of the data.
@@ 3 Perspective effects
We note that the analytically calculated perspective effects given by Equations (10), (11), and (12) agree well with direct numerical calculations with the algorithm described in Appendix B of K.-H. Chae (2025a).
These equations work excellently even for the Proxima - $\alpha$ Cen AB system, the nearest wide binary from the Sun. Our analytic results for this system are expected to show the largest deviation from the direct numerical calculations by P. Kervella et al. (2017). However, the difference is only about 0.07 m/s, totally negligible compared to the real relative velocity of 273 m/s, and far less than 4 m/s, the error of proper motion component velocity, and 32 m/s, the error of radial velocity components. Roughly speaking, the PM motion is given by $\Delta\mu$ in the lowest order, and the perspective effect formulas above give the corrections on the order of $\mathcal{O}(\mu\Delta\theta)$ where $\theta$ denotes the angular separation $\alpha$ and $\delta$. Skipping $\mathcal{O}(\Delta\mu\Delta\theta)$ and $\mathcal{O}(\mu\Delta\theta^{2})$, the next corrections come at the order of $\mathcal{O}(\Delta\mu\Delta\theta^{2}),\mathcal{O}(\mu\Delta\theta^{3})$, which are completely negligible, as we just explained.
@@
is about two orders of magnitude greater than the true value of $\Delta d$, which should be in the order of $s$, the sky-projected separation. See the top-left panel of Figure 2 for a histogram of $\sigma_{\Delta d}/s$. Thus, considering the first term would induce errors that would be much larger than those without it. Lacking accurate values of $\Delta d$, we could estimate the effect of the term proportional to $\Delta d/d$ using realistic MC generated values of $\Delta d$, but it is not warranted in a statistical sense because $\Delta d/d$ can be positive or negative with equal probabilities, see Appendix B of K.-H. Chae (2025a).
@@ 4 Scaling relations
Then $s/r_{M}$ (where $r_{M}$ is given by Equation (18)) and $\tilde{v}$ are dimensionless. According to Newtonian gravity, the mean (or the median) value of $\tilde{v}$ should be roughly constant, independent of $s/r_{M}$ (as long as eccentricity is nearly independent of $s$). However, we clearly see that $\tilde{v}$ deviates from a flat line beyond about $s/r_{M}=0.3$.
@@ 5 Monte Carlo simulations
For the present sample, the two highest acceleration bins with $x_{0}>-8$ satisfy $\delta_{\rm obs-newt}\approx 0$ with $f_{\rm multi}=0$. This means that in an average or median sense, our sample is pure, free of kinematic contaminants. This can be understood from the fact that every binary in our sample satisfies the highest data qualities (e.g., $\sigma_{v_{p}}<50$ m s-1) and stringent selection criteria such as ruwe <1.2<1.2. This will not mean that there are no individual exceptions that evaded our selection criteria, but relatively few exceptions would not affect the median trend. We also emphasize that every star from all our binaries satisfies the same criteria with an equal probability, regardless of their membership, meaning that the exceptional cases would occur regardless of the binary separation and thus cannot bias the median trend we see. Although a close binary of the same mass and separation from the host star can have a relatively larger impact on the relative velocity in the wide binary of a larger separation, it would not affect the median trend significantly as long as its probability is low because a median is immune to a few exceptional outliers.
The purple line in Figure 6 is the AQUAL prediction, as presented in Figure 1 of K.-H. Chae (2024a). We find that the gravitational acceleration boost factor $\gamma_{g}$ (Equation (24)) is $1.61^{+0.37}_{-0.29}$ in the last bin, i.e., at $g_{\rm N}\approx 10^{-11.0}\text{m s}^{-2}$ and $1.26^{+0.12}_{-0.10}$ in the second to last bin, i.e., at $g_{\rm N}\approx 10^{-10.3}\text{m s}^{-2}$. The last bin represents the lowest-acceleration bin probed to date, thanks to our inclusion of wide binaries with $s>30$ kau. Unfortunately, it returns a result with a relatively large uncertainty. Nevertheless, it is intriguing to note that the boost factor $1.61^{+0.37}_{-0.29}$ is consistent with the AQUAL/QUMOND prediction $\approx 1.4-1.5$. Although not shown in the figure, we find $1.32^{+0.12}_{-0.11}$ for $g_{\rm N}\lesssim 10^{-10}\text{m s}^{-2}$ that encompasses the two bins. These boost factors are consistent with previous estimates based on differently defined samples (e.g., K.-H. Chae 2023, 2024a, 2024b, 2025a; X. Hernandez et al. 2024; X. Hernandez & P. Kroupa 2025).
@@
The above results are based on the nominal sample and the 2D velocities corrected for the perspective effects. In Appendix A, we perform the acceleration-plane analysis of our nominal sample without considering the perspective effects, so that the impact of the perspective effects on gravity inference can be seen directly. We see that the perspective effects have a rather minor, though non-negligible, impact on gravity inference. In Appendix B, we consider alternative samples with varied ruwe limits of <1.1<1.1 and <1.3<1.3 instead of the nominal limit <1.2<1.2. We see that there are no significant changes in the gravity inference. In Appendix C, we consider a sample obtained by a stricter criterion for the color-magnitude (CM) relation than the one of Z. Penoyre et al. (2022). We mask the region in the CM diagram that may contain the photometric binary sequence (note however that some in the masked region may well be single stars). It turns out that this cut has a minor impact on gravity inference because those removed are relatively few and therefore cannot affect the median trend significantly. In Appendix D, we consider the sample of binaries with both radial velocities available. Compared to the nominal sample, gravity inference is little changed.